# KiDS-Legacy: angular galaxy clustering from deep surveys with complex selection effects

### v1

@@ 4.3 Reconstructing organised randoms
The basic idea of constructing an OR is to first find the effective sky region containing galaxies from each cluster. Galaxies in this region can be treated as uniformly selected according to the specific synthetic systematics of that cluster. We then generate a uniform random sample in the region according to the mean number density of that cluster and then add up the URs corresponding to all the clusters to get the overall OR. It is desirable to obtain an OR catalogue, i.e. coordinates of points within the observation footprint whose spatial distribution strictly follows the depletion distribution. However, with a finite number of hierarchical clusters, we cannot achieve such a resolution. Instead, we generate pixelized OR maps of the sky. That is, we aim for the OR weight in sky pixels that reflects the selection function for galaxies in that pixel ($P_{\mathrm{keep}}$). When measuring the 2PCF, we use it to calculate the $RR$ term in Eq. (9). In this work, we use the HEALPix scheme to pixelise the sky.
The OR is constructed as follows: after identifying the galaxy clusters in the systematics space via SOM+HC, we first find $N^{i}_{p}$, the number of galaxies in the $i$-th cluster that are in the $p$-th pixel of the sky (note that a pixel can contain galaxies from different clusters). For a given cluster index $i$, $N^{i}_{p}$ should occupy only discrete sub-regions within the footprint that contain galaxies distributed closely in the high-dimensional space. We calculate the effective area of the $p$-th pixel corresponding to the $i$-th cluster as:
@@
Now we add up $A_{p}^{i}$ across pixels to get the total effective area for each cluster:
@@
and the effective surface number density of the $i$-th cluster:
@@
The OR weight $\mathcal{W}_{p}$ is an estimate of $P_{\mathrm{keep}}$, so it is used as the $R$ terms in Eq. (7) to measure the unbiased 2PCF. Note that we use the calligraphic font $\mathcal{W}$ for the OR weight to distinguish from the 2PCF $w$. Figure 1 summarises the SOM+HC method in a flowchart.
The left panel of Figure 6 shows the KiDS-Legacy galaxy number fluctuation colour-coded by galaxy number per pixel. The holes in this map are the masked regions. From the left panel, one can notice an obvious tile-based variable depth in the galaxy map by eye. The right panel shows the organised random weight map constructed with a $30\times 30$ SOM, grouped into 400 hierarchical clusters. Both maps are pixelised with Nside=2048. Visually, it is clear that the OR weight map also shows a tile-based pattern that roughly matches the pattern on the galaxy number map.
Figure 6: Left panel: the galaxy number per pixel on the part of the KiDS-Legacy footprint; right panel: the OR weight on the same region.
Pixel size is another important parameter affecting the performance of the OR reconstruction. If the pixel size is too large, we lose the spatial resolution of the variable depth, resulting in an overestimated 2PCF with residual variable depth correlations. If the pixel size is too small, the organised random weight map will be too sparse, with many unintentionally masked pixels in the organised random map, since the organised random weight will only be non-zero in pixels containing galaxies. Thus the organised random will correlate with the LSS. If we correct this, we will remove the LSS from the 2PCF and underestimate it. Therefore, we need to choose the pixel size carefully to balance the trade-off between resolution and over-correction. For the HEALPix scheme used in this paper, the pixel size is determined by the Nside parameter of the map. Nside can only be a power of 2, and if Nside doubles, the pixel size is halved. Our baseline analysis uses Nside=2048 corresponding to an angular size of 1.7 arcmin. We will validate this choice later.
@@ 5 Validation of the method
In this section, we validate the SOM+HC method with two case studies. In Subsection 5.1, we recover the OR for mock galaxy samples selected by systematics with simple spatial distributions and selection functions, while in Subsection 5.2.2, we recover the OR for mock galaxy samples selected by realistic systematics distributions and data-driven selection functions. We evaluate the method by calculating the bias in the OR-corrected 2PCF.
@@ 5.1 Testing the organised random with toy systematics model
Figure 7: The spatial distribution of the four toy systematics. Systematics A is tile-independent (type-A); Systematics B varies within each tile (type-B); Systematics C1 and C2 uniform in each tile but differ across tiles (type-C). The black curves in the colour bar show the selection function of each systematic.
Figure 8: The flowchart of the toy model validation of the SOM+HC method. Note that the “UR” case ($w(\theta)$ calculated with depleted mock catalogue and mock UR) is not shown in this figure.
@@
  • Systematics A mimics Type A varies smoothly over large angles. We model it as a Gaussian function of Galactic latitude. It mimics large-scale variations such as the Galactic foreground.
@@
  • Systematics C1 varies discretely across tiles but is constant within them, mimicking per-exposure effects such as limiting depth variations (background level, PSF size, etc.) resulting from the use of a step-and-stare observing strategy.
  • Systematics C2 is the same idea as Systematics C1, but with a different realisation.
@@
These four systematics are corresponding to type-A (Systematics A), type-B (Systematics B), and type-C (Systematics C1 and Systematics C2) systematics introduced in Section 3. Their values are normalised to be between 0 and 1. Since SOM+HC does not change the topology of the galaxy distributions in systematics space, this normalisation does not affect the performance of the method.
@@
We define simple selection functions for Systematics A (a trigonometric function modulated by a linear function), Systematics B (a quadratic function), and Systematics C1 (a linear function). Systematics C2 has no selection effect and thus acts as a “distractor” for the SOM. We show the systematics map and the associated selection function in Figure 7. We also assume that the selections between the systematics are independent so that the overall selection function is a multiplication of the selection functions of the individual systematics. To make the selection, we generate a uniform random number between 0 and 1 for each galaxy, and if the random number is less than the selection function $P_{\mathrm{keep}}(\boldsymbol{\theta})$ at its position, the galaxy is kept in the catalogue. After selection, we have a galaxy catalogue with an angular number density of $\sim$1 arcmin-2.
@@
We also generate a “true” OR catalogue by applying the same selection functions to a UR catalogue in the same footprint. If we use it as the $R$ terms for the 2PCF measurement with the depleted mock catalogue, we should get an unbiased $w(\theta)$. Note that in reality, we do not have access to the true OR. In this validation test, we recover the OR with a 30$\times$30 SOM with hexagonal cells and a toroidal topology (so that the horizontal and vertical edges are adjacent), grouped into 200 hierarchical clusters. To check whether the SOM+HC recovers the input selection functions, we plot the relationship between galaxy number contrast (relative difference between galaxy number in each cluster and the mean galaxy number across clusters) and median systematics values of each cluster in Figure 9. The blue curves are the number contrast derived from the input selection rates (the black curves in the colour bars of Figure 7). The median systematics and number contrast for each point are calculated by first sorting the median systematics for all clusters in a realisation, then averaging the median systematics and number contrasts across realisations. Standard errors are also calculated and presented as error bars. The errors of the median systematics are too small to be visible. The galaxy number contrasts of the HCs generally follow the input selection rate, indicating that SOM+HC is able to recover the input selection functions of each systematics. Notably, SOM+HC will not introduce additional correlation for Systematics C2, which does not select galaxies. We also note that the number contrast slightly mismatches the selection functions at the edges. This is probably because the SOM is less effective closer to the boundary in the systematics space, where there are fewer neighbouring galaxies.
@@
Figure 9: The relationship between the galaxy number contrast and the mean systematic value of each hierarchical cluster. The blue curves are the number contrast derived from the input selection rates (the black curves in the colour bars of Figure 7). The average median systematics and number contrast are calculated by averaging the sorted values in each realisation across all the realisations. Standard errors are also calculated and presented as error bars. The errors of the median systematics are too small to be visible.
@@
Figure 10: Top panel: the true OR weights (normalised by their mean) calculated from the total selection function of the toy systematics; middle panel: the recovered OR weights generated by the SOM+HC method. Both panels show only part of the footprint. Note that the holes in the maps are masked regions around “point sources”; bottom panel: the relative difference between the recovered OR weights and the true OR weights.
@@
The recovered OR weight map is pixelised into a HEALPix map with Nside=1024. The true OR weight map, the recovered OR weight map, and the relative difference are shown in Figure 10. Visually, the recovered OR shows the same spatial pattern as the true OR, but it appears more discrete. This is due to a finite number of clusters in the systematics space (in the words, if we had an infinite number of galaxies grouped into an infinite number of clusters, we would recover the smooth OR weights). The relative difference (the bottom panel of Figure 10) is well within ±$\sim\pm 20\%$.
@@
In summary, we have an unselected mock galaxy catalogue and a selected catalogue, plus a uniform random, a true organised random and a recovered organised random. To evaluate the SOM+HC method quantitatively, we measure four $w(\theta)$’s from them. They are summarised in Table 2. In summary, we validate the method by comparing the “Recovered OR” and the “No selection” 2PCFs. The 2PCF is measured in 20 angular bins between 2.5 and 250 arcmin (following DeRose et al., 2022) with the TreeCorr [^12] https://rmjarvis.github.io/TreeCorr/_build/html/index.html package (Jarvis, 2015). To evaluate the covariance matrix and reduce the sample variance, we perform the above validation on 40 GLASS realisations and evaluate the covariance as:
@@
where the subscript $r$ denotes the realisation number, and $\bar{w}(\theta)$ is the average 2PCF over realisations. We also compute a theoretical covariance matrix using the OneCovariance code (Reischke et al., 2024) [^13] https://github.com/rreischke/OneCovariance and find that it is consistent with the “No selection” covariance matrix (see Appendix A for a comparison). We evaluate the bias of the 2PCFs by calculating the $\chi^{2}$ between each $w(\theta)$ and the “no selection” $w(\theta)$:
@@
where $\bar{\boldsymbol{w}}$ is the difference between the 2PCF and the “No selection” 2PCF. Assuming $\chi_{\mathrm{d}}^{2}$ follows a $\chi^{2}$ distribution with degrees-of-freedom equal to the number of angular bins taken into account, we can calculate the corresponding probability-to-exceed (PTE) value. The 2PCF is less biased if the PTE value is closer to 1.
The measured 2PCFs and their relative difference to the “no selection” case are shown in Figure 11. In this work, we focus on linear scales, so we calculate $\chi_{\mathrm{d}}$ only on the angular scale $\theta<\theta_{\mathrm{cut}}$ where the cutting scale $\theta_{\mathrm{cut}}$ is the angular scale corresponding to a physical scale of $8h^{-1}\mathrm{Mpc}$ at the mean redshift of the mock sample ($\bar{z}=0.3$ in our case), which has a value of 42.74 arcmin. For the UR case, $\chi_{\mathrm{d}}=1616$, meaning a huge bias if we use uniform random for the depleted catalogue. For the true OR case, $\chi_{\mathrm{d}}=0.08$ (corresponding to a PTE=0.99999987), meaning that true organised random can give an unbiased $w(\theta)$ as expected. For the “Recovered OR” case, $\chi_{\mathrm{d}}=0.24$ (corresponding to a PTE=0.999992). This means that the OR recovered by the SOM+HC method also gives unbiased $w(\theta)$. Johnston et al. (2021) have shown that SOM+HC can correct a slight variable depth bias in the KiDS-Bright sample. Notably, in this test, SOM+HC can correct $w(\theta)$ that is biased by orders of magnitude. It should also be noted that the wiggle at small scales is due to pixelisation, which changes the angular distance between galaxies as they are effectively moved to the centre of the pixel. This effect particularly affects scales close to the pixel size (3.4 arcmin in our case).
@@
Table 2: The information of the four $w(\theta)$’s measured from the mock sample. The “Label” column indicates the aliases of the $w(\theta)$’s in the following discussion and figures; the “Biased?” column indicates whether the 2PCF is biased; the “Colour” column indicates the colours of the data points shown in the figures. The “No selection” (the first row) and “True OR” (the third row) cases are unbiased benchmarks and the “UR” (the second row) case is a biased measurement. The “Recovered OR” case (the fourth row) is the 2PCF that we use to validate our method.
@@ 5.2 Validation with data-driven systematics
Figure 12: The flowchart of the data-driven test.
@@ 5.2.1 Introduction to the method
To tackle this problem, we apply “data-driven systematics” to the mock catalogue following Johnston et al. (2021). First, we generate a mock galaxy catalogue following a realistic redshift distribution and footprint [^14] If the variable depth does not encode cosmological information, it is not necessary to use a realistic cosmology and redshift distribution to generate the mock.; then we generate the “data-driven mock systematics” by assigning systematics values to each mock galaxy with a nearest neighbour interpolation from the real KiDS systematics. Thus, we get a mock galaxy sample with the same spatial distributions of systematics as the real data. Note that even if the spatial distributions of the systematics are correlated with the LSS, this step removes any such correlation because the mock galaxy catalogue has a different spatial distribution. Next, we train a SOM+HC on the real galaxy catalogue (we call it “SOM+HC+KiDS”) and get an OR weight map $\mathcal{W}_{\mathrm{KiDS}}$ as a proxy of the realistic variable depth (we call it the “data-driven OR weight”); then we select mock galaxies according to it. We define the “OR contrast” as
@@
$P_{\mathrm{keep}}(\boldsymbol{\theta})=\frac{N_{\mathrm{out}}}{N_{\mathrm{in}}}\left[1+\delta_{\mathrm{OR,KiDS}}(\boldsymbol{\theta})\right]\,,$ (20)
where $N_{\mathrm{out}}$ is the desired galaxy number of the depleted mock sample. In this work, we choose $N_{\mathrm{out}}=49,875,861$, the galaxy number in the KiDS-Legacy catalogue. The galaxy number of the input, unselected mock sample $N_{\mathrm{in}}$ is chosen to ensure $P_{\mathrm{keep}}\leq 1$. The selection procedure is the same as described in Subsection 5.1. For each galaxy we generate a uniform random number between 0 and 1 and compare it with $P_{\mathrm{keep}}$ associated with the pixel containing the galaxy. If the random number is less than $P_{\mathrm{keep}}$, we keep the galaxy in the sample; otherwise, we discard it.
The assignment of mock systematics with nearest-neighbour interpolation ensures that the spatial distribution of the mock systematics is realistic. If SOM+HC recovers an unbiased OR weight on the real data, the data-driven OR weight should represent the realistic variable depth. When applied as a selection function to the mock data, it should produce the variable depth caused by the realistic selection function of the mock systematics. We run SOM+HC on the post-selected mock catalogue with the mock systematics to test this consistency. Ideally, the resulting OR weight should match the imported selections, i.e. the data-driven OR weight, and we expect to measure an unbiased 2PCF from the selected mock catalogue corrected with the recovered mock OR. Figure 12 shows the flowchart of the data-driven systematics test.
@@
Several configuration parameters affect the accuracy of the recovered OR weight. In particular, the SOM+HC procedure depends on $NC$, the number of hierarchical clusters, and Nside, the OR weight resolution. If $NC$ and/or Nside is too low, SOM+HC will fail to resolve systematic clustering in systematic/real space, resulting in uncorrected variable depth. If $NC$ is too high, then SOM+HC will start to resolve LSS-induced clustering in systematics space. If Nside is so high that significant amounts of pixels get no galaxy, then the OR weight will follow the LSS. In this section, we aim to perform data-driven systematics tests to find the optimal {$NC$, Nside}. Note that the data-driven systematics test depends on these parameters for both data-driven and recovered mock OR. In the following discussion, we use {$NC_{\mathrm{KiDS}}$, $\texttt{Nside}_{\mathrm{KiDS}}$} to denote the setup for the data-driven OR, and {$NC_{\mathrm{rec}}$, $\texttt{Nside}_{\mathrm{rec}}$} for the recovered mock OR. To avoid endless tests for different parameter value combinations, we assume that the same SOM+HC setups on real and mock data, with the same systematics choices, will yield equally good or bad performance on both real and mock data. With this assumption in mind, we can evaluate the performance of a $\{NC_{\mathrm{KiDS}},\texttt{Nside}_{\mathrm{KiDS}}\}$ combination based on the recovered OR 2PCF with $NC_{\mathrm{KiDS}}=NC_{\mathrm{rec}},\texttt{Nside}_{\mathrm{KiDS}}=\texttt{Nside}_{\mathrm{rec}}$:
@@
    If the recovered OR 2PCF is significantly lower than the “no selection” case, the $\{NC_{\mathrm{KiDS}},\texttt{Nside}_{\mathrm{KiDS}}\}$ option is likely to overcorrect the variable depth by removing the LSS;
@@
    If the recovered OR 2PCF agrees with the “No selection” case, it is possible that the choice is optimal. It is also possible that the data-driven OR underestimates the variable depth and induces too soft selections, which can be corrected by the same $\{NC_{\mathrm{rec}},\texttt{Nside}_{\mathrm{rec}}\}$ combination, while a realistic variable depth actually requires higher $\{NC_{\mathrm{KiDS}},\texttt{Nside}_{\mathrm{KiDS}}\}$. For the second case, one can think of an extreme case: $NC_{\mathrm{KiDS}}=NC_{\mathrm{rec}}=1$. This case is the same as “no selection” and will give an unbiased 2PCF, but the number of clusters is clearly too low for the real data with complex selection effects.

### v2

@@ 4.3 Reconstructing organised randoms
The basic idea of constructing an OR is to first find the effective sky region containing galaxies from each cluster. Galaxies in this region can be treated as uniformly selected according to the specific synthetic systematics of that cluster. We then generate a uniform random sample in the region according to the mean number density of that cluster and then add up the URs corresponding to all the clusters to get the overall OR. It is desirable to obtain an OR catalogue (i.e. coordinates of points within the observation footprint whose spatial distribution strictly follows the depletion distribution). However, with a finite number of hierarchical clusters, we cannot achieve such a resolution. Instead, we generate pixelised OR maps of the sky. That is, we aim for the OR weight in sky pixels that reflects the selection function for galaxies in that pixel ($P_{\mathrm{keep}}$). When measuring the 2PCF, we use it to calculate the $RR$ term in Eq. (9). In this work, we use the HEALPix scheme to pixelise the sky.
The OR is constructed as follows: after identifying the galaxy clusters in the systematics space via SOM+HC, we first find $N^{i}_{p}$, the number of galaxies in the $i$-th cluster that are in the $p$-th pixel of the sky; we note that a pixel can contain galaxies from different clusters. For a given cluster index $i$, $N^{i}_{p}$ should occupy only discrete sub-regions within the footprint that contain galaxies distributed closely in the high-dimensional space. We calculate the effective area of the $p$-th pixel corresponding to the $i$-th cluster as
@@
Now we add up $A_{p}^{i}$ across pixels to get the total effective area for each cluster,
@@
and the effective surface number density of the $i$-th cluster,
@@
The OR weight $\mathcal{W}_{p}$ is an estimate of $P_{\mathrm{keep}}$, so it is used as the $R$ terms in Eq. (7) to measure the unbiased 2PCF. We note that we use the calligraphic font $\mathcal{W}$ for the OR weight to distinguish from the 2PCF $w$. Figure 4 summarises the SOM+HC method in a flowchart.
The left panel of Fig. 6 shows the KiDS-Legacy galaxy number fluctuation colour-coded by galaxy number per pixel. The holes in this map are the masked regions. From the left panel, one can notice an obvious tile-based variable depth in the galaxy map by eye. The right panel shows the organised random weight map constructed with a $30\times 30$ SOM, grouped into 400 hierarchical clusters. Both maps are pixelised with Nside=2048. Visually, it is clear that the OR weight map also shows a tile-based pattern that roughly matches the pattern on the galaxy number map.
Figure 6: Left panel: Galaxy number per pixel on a subregion of the KiDS-Legacy footprint; right panel: OR weight on the same region.
Pixel size is another important parameter affecting the performance of the OR reconstruction. If the pixel size is too large, we lose the spatial resolution of the variable depth, resulting in an overestimated 2PCF with residual variable depth correlations. If the pixel size is too small, the organised random weight map will be too sparse, with many unintentionally masked pixels in the organised random map, since the organised random weight will only be non-zero in pixels containing galaxies. Thus the organised random will correlate with the LSS. If we correct this, we will remove the LSS from the 2PCF and under-estimate it. Therefore, we need to choose the pixel size carefully to balance the trade-off between resolution and over-correction. For the HEALPix scheme used in this paper, the pixel size is determined by the Nside parameter of the map. Nside can only be a power of 2, and if Nside doubles, the pixel size is halved. Our baseline analysis uses Nside=2048 corresponding to an angular size of 1.7 arcmin. We validate this choice below.
@@ 5 Validation of the method
In this section we validate the SOM+HC method with two case studies. In Sect. 5.1, we recover the OR for mock galaxy samples selected by systematics with simple spatial distributions and selection functions, while in Sect. 5.2.2, we recover the OR for mock galaxy samples selected by realistic systematics distributions and data-driven selection functions. We evaluate the method by calculating the bias in the OR-corrected 2PCF.
@@ 5.1 Testing the organised random with toy systematics model
Figure 7: Left panels: Spatial distribution of the four toy systematics. Systematics A1 and A2 are uniform in each tile, but differ across tiles (type A); Systematics B varies within each tile (type B); Systematics C is tile-independent (type C); Right panels: Black curves in the colour bar show the selection function of each systematic and the shaded region is the normalised distribution of the systematics. The numbers on the right show the selection rate values.
Figure 8: Flowchart of the toy model validation of the SOM+HC method. We note that the UR case ($w(\theta)$ calculated with depleted mock catalogue and mock UR) is not shown in this figure.
@@
  • Systematics A1 varies discretely across tiles but is constant within them, mimicking per-exposure effects such as limiting depth variations (e.g. background level, PSF size) resulting from the use of a step-and-stare observing strategy.
  • Systematics A2 is the same idea as Systematics A1, but with a different realisation.
@@
  • Systematics C varies smoothly over large angles. We model it as a Gaussian function of Galactic latitude. It mimics large-scale variations such as the Galactic foreground.
@@
These four systematics are corresponding to type A (Systematics A1 and Systematics A2), type B (Systematics B), and type C (Systematics C) systematics introduced in Sect. 3. Their values are normalised to be between 0 and 1. Since SOM+HC does not change the topology of the galaxy distributions in systematics space, this normalisation does not affect the performance of the method.
@@
We define simple selection functions for Systematics A1 (a linear function), Systematics B (a quadratic function), and Systematics C (a trigonometric function modulated by a linear function). Systematics A2 has no selection effect and thus acts as a “distractor” for the SOM. We show the systematics map and the associated selection function in Fig. 7. We also assume that the selections between the systematics are independent so that the overall selection function is a multiplication of the selection functions of the individual systematics. To make the selection, we generate a uniform random number between 0 and 1 for each galaxy, and if the random number is smaller than the selection function $P_{\mathrm{keep}}(\boldsymbol{\theta})$ at its position, the galaxy is kept in the catalogue. After selection, we have a galaxy catalogue with an angular number density of $\sim$1 arcmin-2.
@@
We also generate a “true” OR catalogue by applying the same selection functions to a UR catalogue in the same footprint. If we use it as the $R$ terms for the 2PCF measurement with the depleted mock catalogue, we should get an unbiased $w(\theta)$. We note that in reality, we do not have access to the true OR. In this validation test, we recover the OR with a 30$\times$30 SOM with hexagonal cells and a toroidal topology (so that the horizontal and vertical edges are adjacent), grouped into 200 hierarchical clusters. To check whether the SOM+HC recovers the input selection functions, we plot the relationship between galaxy number contrast (relative difference between galaxy number in each cluster and the mean galaxy number across clusters) and median systematics values of each cluster in Fig. 9. The blue curves are the number contrast derived from the input selection rates (the black curves in the colour bars of Fig. 7). The median systematics and number contrast for each point are calculated by first sorting the median systematics for all clusters in a realisation, then averaging the median systematics and number contrasts across realisations. Standard errors are also calculated and presented as error bars. The errors of the median systematics are too small to be visible. The galaxy number contrasts of the HCs generally follow the input selection rate, indicating that SOM+HC is able to recover the input selection functions of each systematics. Notably, SOM+HC will not introduce additional correlation for Systematics A2, which does not select galaxies. We also note that the number contrast slightly mismatches the selection functions at the edges. This is probably because the SOM is less effective closer to the boundary in the systematics space, where there are fewer neighbouring galaxies.
@@
Figure 9: Relationship between the galaxy number contrast and the mean systematic value of each hierarchical cluster. The blue curves are the number contrast derived from the input selection rates (black curves in the colour bars of Fig. 7). The average median systematics and number contrast are calculated by averaging the sorted values in each realisation across all the realisations. The standard errors are also calculated and presented as error bars. The errors of the median systematics are too small to be visible.
@@
Figure 10: Top panel: True OR weights (normalised by their mean) calculated from the total selection function of the toy systematics; middle panel: Recovered OR weights generated by the SOM+HC method. Both panels show only part of the footprint. We note that the holes in the maps are masked regions around point sources; bottom panel: Relative difference between the recovered OR weights and the true OR weights.
@@
The recovered OR weight map is pixelised into a HEALPix map with Nside=1024  [^13] The Nside used here is lower than the optimal Nside for the real data because we generate fewer mock galaxies in this test, and Nside=1024 ensures that each pixel has $\sim$10 galaxies.. The true OR weight map, the recovered OR weight map, and the relative difference are shown in Fig. 10. Visually, the recovered OR shows the same spatial pattern as the true OR, but it appears more discrete. This is due to a finite number of clusters in the systematics space (in other words, if we had an infinite number of galaxies grouped into an infinite number of clusters, we would recover the smooth OR weights). The relative difference (the bottom panel of Fig. 10) is well within ±$\sim\pm 20\%$.
@@
In summary, we have an unselected mock galaxy catalogue and a selected catalogue, plus a uniform random, a true organised random and a recovered organised random. To evaluate the SOM+HC method quantitatively, we measure four $w(\theta)$’s from them. They are summarised in Table 16. In summary, we validate the method by comparing the “Recovered OR” and the “No selection” 2PCFs. The 2PCF is measured in 20 angular bins between 2.5 and 250 arcmin (following DeRose et al., 2022) with the TreeCorr [^14] https://rmjarvis.github.io/TreeCorr/_build/html/index.html package (Jarvis, 2015). To evaluate the covariance matrix and reduce the sample variance, we perform the above validation on 40 GLASS realisations and evaluate the covariance as
@@
where the subscript $r$ denotes the realisation number, and $\bar{w}(\theta)$ is the average 2PCF over realisations. We also compute a theoretical covariance matrix using the OneCovariance code (Reischke et al., 2024) [^15] https://github.com/rreischke/OneCovariance and find that it is consistent with the “No selection” covariance matrix (see Appendix A for a comparison). We evaluate the bias of the 2PCFs by calculating the $\chi^{2}$ between each $w(\theta)$ and the “No selection” $w(\theta)$
@@
where $\Delta\bar{\boldsymbol{w}}$ is the difference between the 2PCF and the “No selection” 2PCF. Assuming $\chi_{\mathrm{d}}^{2}$ follows a $\chi^{2}$ distribution with degrees-of-freedom equal to the number of angular bins taken into account, we can calculate the corresponding probability-to-exceed (PTE) value. The 2PCF is less biased if the PTE value is closer to 1.
The measured 2PCFs and their relative difference to the “No selection” case are shown in Fig. 11. In this work, we focus on linear scales, so we calculate $\chi_{\mathrm{d}}$ only on the angular scale $\theta<\theta_{\mathrm{cut}}$ where the cutting scale $\theta_{\mathrm{cut}}$ is the angular scale corresponding to a physical scale of $8h^{-1}\mathrm{Mpc}$ at the mean redshift of the mock sample ($\bar{z}=0.3$ in our case), which has a value of 42.74 arcmin. For the UR case, $\chi_{\mathrm{d}}=1616$, meaning a huge bias if we use uniform random for the depleted catalogue. For the true OR case, $\chi_{\mathrm{d}}=0.08$ (corresponding to a PTE=0.99999987), meaning that true organised random can give an unbiased $w(\theta)$ as expected. For the “Recovered OR” case, $\chi_{\mathrm{d}}=0.24$ (corresponding to a PTE=0.999992). This means that the OR recovered by the SOM+HC method also gives unbiased $w(\theta)$. Johnston et al. (2021) have shown that SOM+HC can correct a slight variable depth bias in the KiDS-Bright sample. Notably, in this test, SOM+HC can correct $w(\theta)$ that is biased by orders of magnitude. It should also be noted that the wiggle at small scales is due to pixelisation, which changes the angular distance between galaxies as they are effectively moved to the centre of the pixel. This effect particularly affects scales close to the pixel size (3.4 arcmin in our case).
Table 2: Information for the four $w(\theta)$ measured from the mock sample.
@@
 [^16] The ‘Label’ column indicates the aliases of the $w(\theta)$ values in the following discussion and figures; the ‘Biased?’ column indicates whether the 2PCF is biased; the ‘Colour’ column indicates the colours of the data points shown in the figures. The ‘No selection’ (first row) and ‘True OR’ (third row) cases are unbiased benchmarks and the ‘UR’ (second row) case is a biased measurement. The ‘Recovered OR’ case (fourth row) is the 2PCF that we used to validate our method.
@@ 5.2 Validation with data-driven systematics
Figure 12: Flowchart of the data-driven test.
@@ 5.2.1 Introduction to the method
To tackle this problem, we apply “data-driven systematics” to the mock catalogue following Johnston et al. (2021). First, we generate a mock galaxy catalogue following the fiducial cosmology given in Sect. 1 with a galaxy bias $b=1$. The galaxy sample follows the same calibrated redshift distribution of the KiDS-Legacy sample  [^17] If the variable depth does not encode cosmological information, it is not necessary to use a realistic cosmology and redshift distribution to generate the mock. and is generated within the KiDS-Legacy footprint; then we generate the “data-driven mock systematics” by assigning systematics values to each mock galaxy with a nearest neighbour interpolation from the real KiDS systematics. Thus, we get a mock galaxy sample with the same spatial distributions of systematics as the real data. We note that even if the spatial distributions of the systematics are correlated with the LSS, this step removes any such correlation because the mock galaxy catalogue has a different spatial distribution. Next, we train a SOM+HC on the real galaxy catalogue (we call it “SOM+HC+KiDS”) and get an OR weight map $\mathcal{W}_{\mathrm{KiDS}}$ as a proxy of the realistic variable depth (we call it the “data-driven OR weight”); then we select mock galaxies according to it. We define the “OR contrast” as
@@
$P_{\mathrm{keep}}(\boldsymbol{\theta})=\frac{N_{\mathrm{out}}}{N_{\mathrm{in}}}\left[1+\delta_{\mathrm{OR,KiDS}}(\boldsymbol{\theta})\right]\,.$ (20)
Here $N_{\mathrm{out}}$ is the desired galaxy number of the depleted mock sample. In this work, we choose $N_{\mathrm{out}}=49,875,861$, the galaxy number in the KiDS-Legacy catalogue. The galaxy number of the input, unselected mock sample $N_{\mathrm{in}}$ is chosen to ensure $P_{\mathrm{keep}}\leq 1$. The selection procedure is the same as described in Sect. 5.1. For each galaxy we generate a uniform random number between 0 and 1 and compare it with $P_{\mathrm{keep}}$ associated with the pixel containing the galaxy. If the random number is less than $P_{\mathrm{keep}}$, we keep the galaxy in the sample; otherwise, we discard it.
The assignment of mock systematics with nearest-neighbour interpolation ensures that the spatial distribution of the mock systematics is realistic. If SOM+HC recovers an unbiased OR weight on the real data, the data-driven OR weight should represent the realistic variable depth. When applied as a selection function to the mock data, it should produce the variable depth caused by the realistic selection function of the mock systematics. We run SOM+HC on the post-selected mock catalogue with the mock systematics to test this consistency. Ideally, the resulting OR weight should match the imported selections (i.e. the data-driven OR weight), and we expect to measure an unbiased 2PCF from the selected mock catalogue corrected with the recovered mock OR. Figure 12 shows the flowchart of the data-driven systematics test.
@@
Several configuration parameters affect the accuracy of the recovered OR weight. In particular, the SOM+HC procedure depends on $NC$, the number of hierarchical clusters, and Nside, the OR weight resolution. If $NC$ and/or Nside is too low, SOM+HC will fail to resolve systematic clustering in systematic/real space, resulting in uncorrected variable depth. If $NC$ is too high, then SOM+HC will start to resolve LSS-induced clustering in systematics space. If Nside is so high that significant amounts of pixels get no galaxy, then the OR weight will follow the LSS. In this section, we aim to perform data-driven systematics tests to find the optimal {$NC$, Nside}. We note that the data-driven systematics test depends on these parameters for both data-driven and recovered mock OR. In the following discussion, we use {$NC_{\mathrm{KiDS}}$, $\texttt{Nside}_{\mathrm{KiDS}}$} to denote the set-up for the data-driven OR, and {$NC_{\mathrm{rec}}$, $\texttt{Nside}_{\mathrm{rec}}$} for the recovered mock OR. To avoid endless tests for different parameter value combinations, we assume that the same SOM+HC set-ups on real and mock data, with the same systematics choices, will yield equally good or bad performance on both real and mock data. With this assumption in mind, we can evaluate the performance of a $\{NC_{\mathrm{KiDS}},\texttt{Nside}_{\mathrm{KiDS}}\}$ combination based on the recovered OR 2PCF with $NC_{\mathrm{KiDS}}=NC_{\mathrm{rec}},\texttt{Nside}_{\mathrm{KiDS}}=\texttt{Nside}_{\mathrm{rec}}$:
@@
    If the recovered OR 2PCF is significantly lower than the “No selection” case, the $\{NC_{\mathrm{KiDS}},\texttt{Nside}_{\mathrm{KiDS}}\}$ option is likely to over-correct the variable depth by removing the LSS;
@@
    If the recovered OR 2PCF agrees with the “No selection” case, it is possible that the choice is optimal. It is also possible that the data-driven OR under-estimates the variable depth and induces too soft selections, which can be corrected by the same $\{NC_{\mathrm{rec}},\texttt{Nside}_{\mathrm{rec}}\}$ combination, while a realistic variable depth actually requires higher $\{NC_{\mathrm{KiDS}},\texttt{Nside}_{\mathrm{KiDS}}\}$. For the second case, one can think of an extreme case: $NC_{\mathrm{KiDS}}=NC_{\mathrm{rec}}=1$. This case is the same as “No selection” and will give an unbiased 2PCF, but the number of clusters is clearly too low for the real data with complex selection effects.